The salary of a person was increased by 15%. If his salary is now is Rs 9936, what was his initial salary?
Answer by Maths68(1474)

New salary after increase = Rs. 9936
Let x be the Initial salary
Initial salary + 15% of initial salary = New Salary
x+15%(x)=9936
x+(15/100)(x)=9936
x+(3/20)(x)=9936
x+3x/20=9936
(20x+3x)/20=9936
23x=9936*20
23x=198720
x=198720/23
x=8640
Rs. 8640 was his initial salary.